QUESTION 4

A car is purchased with a 48-month, 9% nominal loan with an initial balance of $15,000. What is the balance due halfway through the 4 years.

$7,489

$9,123

$6,094

$8,176

QUESTION 3

Find EUAW of the following project. Use 8 years for the analysis and no renewal, along with i= 10%. The initial cost is $2,800; EUAB is $900 and its useful life is 5 years.

$334.51

$497.11

$850.23

$114.67
Answer #1

4) Monthly payment is 15000 (A/P, 9%/12, 48). Balance due halfway = monthly payment x (P/A, 9%/12, 48 - 24)

= 15000 (A/P, 9%/12, 48) * (P/A, 9%/12, 24)

= 15000 * 0.02489 * 21.88915

= 8173

Select 8176.

5) EUAW = -2800(A/P, 10%, 8) + 900(P/A, 10%, 5)(A/P, 10%, 8)

= -2800*0.18744 + 900*3.79079*0.18744

= 114.67